OpenStudy (anonymous):

Express each quotient in the simplest form. (10xy^2)/(3z) / (5xy)/(6z^3)

OpenStudy (anonymous):

@One098

OpenStudy (leonardo0430):

this might help http://openstudy.com/study#/updates/53b877c7e4b09f140ca65782

OpenStudy (jhannybean):

\[\huge \dfrac{\frac{10xy^2}{3z}}{\frac{5xy}{6z^3}}\]The second fraction, in this case, will be flipped and multiplied to the first. Remember: \(\large \dfrac{\frac{a}{b}}{\frac{c}{d}} = \frac{a}{b} \cdot \frac{d}{c}\) . You can applythe same rule here.

OpenStudy (jhannybean):

\[\large \frac{10xy^2}{3z} \cdot \frac{6z^3}{5xy}\]Cross cancel any like terms. \[\large \frac{10\color{red}{xy^2}}{3\color{blue}z} \cdot \frac{6\color{blue}{z^3}}{5\color{red}{xy}}\]

OpenStudy (anonymous):

no\[4yx^2\] right?